OpenStudy (anonymous):

Find the perimeter of a triangle with vertices (2,4) (-1,3) and (2,-3)

OpenStudy (accessdenied):

\[ D_{distance} = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2} \] Use this for all three sides and add together for perimeter

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no i think i messed it up again

OpenStudy (accessdenied):

i didnt get sqrt(58) -- -3 - 3 = -6 2 - -1 = 2 + 1 = 3 (-6)^2 + 3^2 = 36 + 9 = 45

OpenStudy (accessdenied):

i got sqrt(10) and 7, but sqrt(45) = 3sqrt(5) for second and third point's distance

myininaya (myininaya):